# U-centrality: A Network Centrality Measure Based on Minimum Energy Control for Laplacian Dynamics

### v1

@@ I Introduction
We denote the set of real numbers by $\mathbb{R}$ and the vector space of $n$-dimensional real-valued column vectors by $\mathbb{R}^{n}$. We use bold lower-case letters to denote column vectors. We use $I$ to denote the identity matrix of a known underlying dimension $n$. We use $\mathbf{1}$, $\boldsymbol{0}$, and $\mathbf{e}_{i}$ to denote the all-one vector, the all-zero vector, and the vector with 1 in the $i$th coordinate and 0’s elsewhere in $\mathbb{R}^{n}$, respectively. We use $[n]$ to denote $\{1,\ldots,n\}$. We use $\|\cdot\|_{2}$ to denote the standard Euclidean norm. We use $\mathcal{R}(A)$ to denote the column range of matrix $A$. For an undirected graph $\mathcal{G}=([n],\mathcal{E})$ with $n$ vertices and a set of edges $\mathcal{E}$, define $\mathcal{A}$ to be the adjacency matrix of $\mathcal{G}$ with $\mathcal{A}_{ij}=1$ if $(i,j)\in\mathcal{E}$ and $\mathcal{A}_{ij}=0$, otherwise. We say $h(t)=\mathcal{O}(g(t))$ if there exists a positive real number $M$ and a real number $t_{0}$ such that $|h(t)|\leq M|g(t)|$} for all $t\geq t_{0}$.
@@ II-C2 Control a single node $i$
Observe that when all nodes are controlled, the minimum energy control input steers the state to the consensus state $\frac{c}{n}\mathbf{1}$. When only node $i$ is controlled, although the minimum energy $E_{i}$ remains constant for all $i$, the terminal points $\boldsymbol{x}_{fi}$ differ for each $i$. Therefore, it is natural to consider a node that leads to a better homogeneity, a more central node. More precisely, we can consider the distance $\|\boldsymbol{x}_{fi}-\frac{c}{n}\mathbf{1}\|_{2}$ as a centrality measure. In other words, given an opinion dynamics model over the network, considering the minimum energy control to drive the state from $\boldsymbol{0}$ to achieve an average state $\frac{c}{n}$ within time $t_{f}$, we rank the agents on the network by the distance between the respective terminal states and $\frac{c}{n}\mathbf{1}$ (that is the homogeneous state derived by controlling all agents). Essentially, this centrality measure characterizes the node’s ability to unify the agents state.
@@ III The Main Results
Since 0 is an eigenvalue of $L$ and its algebraic (and hence, geometric) multiplicity is $1$ for a connected graph [23] and $L$ is positive semidefinite, we can diagonalize $L$ as
@@
First, it should be noted that $\left\|\boldsymbol{x}_{fi}-\frac{c}{n}\mathbf{1}\right\|_{2}$ decreases at the rate of $\mathcal{O}\left(\frac{1}{t_{f}}\right)$ as $t_{f}$ increases and when $t_{f}\rightarrow\infty$, $\boldsymbol{x}_{fi}\rightarrow\frac{c}{n}\mathbf{1}$. This makes sense as the larger $t_{f}$ is, the more time agents have to interact through the network so that $\boldsymbol{x}_{fi}$ can be closer to the consensus $\frac{c}{n}\mathbf{1}$. Second, it is more interesting to connect $\|L^{\dagger}\mathbf{e}_{i}\|_{2}$ with node $i$’s position in the graph. In the existing literature, $L^{\dagger}$ is particularly useful to define the resistance distance between nodes in a network [5]. Consider a network in which the edges represent resistors and the nodes are junctions between resistors. Each edge has conductance $1$. Given nodes $i$ and $j$, the resistance distance $R_{ij}$ between nodes $i$ and $j$ is defined as the effective resistance between nodes $i$ and $j$, which is equal to the potential difference between nodes $i$ and $j$ when a current source of 1 Amp is placed between $i$ (input) and $j$ (output). In this case, it has been derived in [5] that
@@ III-A U-centrality of Trees for $t_{f}\gg 0$
Notice that the first and the last term in the right-hand side of (24) are common for all $i\in[n]$. The second term in (24) is the empirical “variance” of the pairwise distance between the node $i$ and the other nodes in the graph, i.e., the empirical variance of $d_{i1},\ldots,d_{in}$, and is small if node $i$ has distances to other nodes that are relatively uniform. Interpreting the third term in (24) is more complicated. Notice that $\frac{D_{i}}{n}-d_{ij}$ is large (in particular positive) if node $j$ is close to $i$, and is small (in particular negative) if node $j$ is far from $i$. Thus, if node $i$ is central, for node $j$ that is close to $i$, node $j$ is also central, so $D_{j}$ is small and $\frac{D_{i}}{n}-d_{ij}$ is large; for node $j$ that is far from $i$, node $j$ is peripheral, so $D_{j}$ is large and $\frac{D_{i}}{n}-d_{ij}$ is small. Therefore, a more central node has a smaller third term. In sum, if node $i$ is more central and other nodes are more ”evenly” positioned relative to it, then $\|L^{\dagger}\mathbf{e}_{i}\|_{2}$ is smaller.
@@ IV Numerical Experiments
  • Degree centrality: The centrality of node $i$ is simply its degree (i.e., the number of nodes connected to $i$). The central nodes are the nodes with the highest degree.

### v2

@@ I Introduction
We denote the set of real numbers by $\mathbb{R}$ and the vector space of $n$-dimensional real-valued column vectors by $\mathbb{R}^{n}$. We use bold lower-case letters to denote column vectors. We use II to denote the identity matrix of a known underlying dimension $n$. We use $\mathbf{1}$, $\boldsymbol{0}$, and $\mathbf{e}_{i}$ to denote the all-one vector, the all-zero vector, and the vector with 1 in the $i$th coordinate and 0’s elsewhere in $\mathbb{R}^{n}$, respectively. We use $[n]$ to denote $\{1,\ldots,n\}$. We use $\|\cdot\|_{2}$ to denote the standard Euclidean norm. We use $\mathcal{R}(A)$ to denote the column range of matrix AA. For an undirected graph $\mathcal{G}=([n],\mathcal{E})$ with $n$ vertices and a set of edges $\mathcal{E}$, define $\mathcal{A}$ to be the adjacency matrix of $\mathcal{G}$ with $\mathcal{A}_{ij}=1$ if $(i,j)\in\mathcal{E}$ and $\mathcal{A}_{ij}=0$, otherwise. We say $h(t)=\mathcal{O}(g(t))$ if there exists a positive real number $M$ and a real number $t_{0}$ such that $|h(t)|\leq M|g(t)|$} for all $t\geq t_{0}$.
@@ II-C2 Control a single node $i$
Observe that when all nodes are controlled, the minimum energy control input steers the state to the consensus state $\frac{c}{n}\mathbf{1}$. When only node $i$ is controlled, although the minimum energy $E_{i}$ remains constant for all $i$, the terminal points $\boldsymbol{x}_{fi}$ differ for each ii. Therefore, it is natural to consider a node that leads to a better homogeneity, a more central node. More precisely, we can consider the distance $\|\boldsymbol{x}_{fi}-\frac{c}{n}\mathbf{1}\|_{2}$ as a centrality measure. In other words, given an opinion dynamics model over the network, considering the minimum energy control to drive the state from $\boldsymbol{0}$ to achieve an average state $\frac{c}{n}$ within time $t_{f}$, we rank the agents on the network by the distance between the respective terminal states and $\frac{c}{n}\mathbf{1}$ (that is the homogeneous state derived by controlling all agents). Essentially, this centrality measure characterizes the node’s ability to unify the agents state.
@@ III The Main Results
Since 0 is an eigenvalue of $L$ and its algebraic (and hence, geometric) multiplicity is 11 for a connected graph [23] and $L$ is positive semidefinite, we can diagonalize $L$ as
@@
First, it should be noted that $\left\|\boldsymbol{x}_{fi}-\frac{c}{n}\mathbf{1}\right\|_{2}$ decreases at the rate of $\mathcal{O}\left(\frac{1}{t_{f}}\right)$ as $t_{f}$ increases and when $t_{f}\rightarrow\infty$, $\boldsymbol{x}_{fi}\rightarrow\frac{c}{n}\mathbf{1}$. This makes sense as the larger $t_{f}$ is, the more time agents have to interact through the network so that $\boldsymbol{x}_{fi}$ can be closer to the consensus $\frac{c}{n}\mathbf{1}$. Second, it is more interesting to connect $\|L^{\dagger}\mathbf{e}_{i}\|_{2}$ with node $i$’s position in the graph. In the existing literature, $L^{\dagger}$ is particularly useful to define the resistance distance between nodes in a network [5]. Consider a network in which the edges represent resistors and the nodes are junctions between resistors. Each edge has conductance 11. Given nodes $i$ and $j$, the resistance distance $R_{ij}$ between nodes $i$ and $j$ is defined as the effective resistance between nodes $i$ and $j$, which is equal to the potential difference between nodes $i$ and $j$ when a current source of 1 Amp is placed between $i$ (input) and $j$ (output). In this case, it has been derived in [5] that
@@ III-A U-centrality of Trees for $t_{f}\gg 0$
Notice that the first and the last term in the right-hand side of (24) are common for all $i\in[n]$. The second term in (24) is the empirical “variance” of the pairwise distance between the node $i$ and the other nodes in the graph, i.e., the empirical variance of $d_{i1},\ldots,d_{in}$, and is small if node $i$ has distances to other nodes that are relatively uniform. Interpreting the third term in (24) is more complicated. Notice that $\frac{D_{i}}{n}-d_{ij}$ is large (in particular positive) if node $j$ is close to $i$, and is small (in particular negative) if node $j$ is far from ii. Thus, if node $i$ is central, for node $j$ that is close to $i$, node $j$ is also central, so $D_{j}$ is small and $\frac{D_{i}}{n}-d_{ij}$ is large; for node $j$ that is far from $i$, node $j$ is peripheral, so $D_{j}$ is large and $\frac{D_{i}}{n}-d_{ij}$ is small. Therefore, a more central node has a smaller third term. In sum, if node $i$ is more central and other nodes are more ”evenly” positioned relative to it, then $\|L^{\dagger}\mathbf{e}_{i}\|_{2}$ is smaller.
@@ IV Numerical Experiments
  • Degree centrality: The centrality of node $i$ is simply its degree (i.e., the number of nodes connected to ii). The central nodes are the nodes with the highest degree.